OpenStudy (anonymous):

1.- a ball is hurled straight up at a speed of 15.0m/s, leaving the hand of the thrower 3.00 m above ground. (a) compute the times and the ball's speed when it passes an observer sitting at a window in line with the trow 10.0 meters above the point of release. (b) What is the total distance traveled by the ball just as it passes the window the second time? 2.- Michael drops a penny into a river from a bridge 44 m above the water. Another penny is thrown vertically down by his friend Jeremy 1.00 second after . Both stones strike the water at the same time. What is the initial speed of the se

OpenStudy (anonymous):

(b) first calculate the distance to top of the flight u = 15, v = 0, g = -9.8, s = ? v^2 = u^2 + 2gs 0 = 225 - 19.6s s = 225/19.6 = 11.48 m total distance travelled = 11.48 + (11.48-10) = 12.96 m